two angles are supplementary angles. The measure of the larger angle is 5 degrees more than 3 times the measure of the smaller angle. find the measure of the two angles.
Answer by nyc_function(2741)

Let x = bigger angle

Let y = smaller angle

x + y = 180

x = 3y + 5

I will solve by substitution.

3y + 5 + y = 180


4y + 5 = 180


4y = 180 - 5


4y = 175


y = 175/4


y = 43.75 degrees.


The smaller angle is 43.75 degrees.


The larger angle is 180 - 43.75 or 136.25 degrees.